OpenStudy (anonymous):

A company that packages cold cuts needs to transport the packaged meat to the supermarket. It will cost $1,200 to rent a truck, and it will cost an additional $100 for every 50 pounds transported. Which equation represents the transport cost for every 50 pounds of cold cuts transported?

OpenStudy (anonymous):

y = 1,200 − 100x y = 1,200 + 100x y = 1,200x + 100 y = 100x − 1,200

OpenStudy (texaschic101):

x = # of 50 lbs of cold cuts y = 1200 + 100x -- (1200 is the truck rental + 100x)

OpenStudy (anonymous):

thanks^